ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 86, 1994

Alfandary, G. On graphs related to conjugacy classes of groups  211
Aljadeff, E. On the surjectivity of some trace maps  221

Altshuler, A., J. Bokowski and P. Schuchert Spatial polyhedra without diagonals

373

Barnette, D. W. A construction of 3-connected graphs  397

Berenstein, C. and D. Pascuas Morera and mean-value type theorems in the hyperbolic disk

61

Bhattacharjee, M. The probability of generating certain profinite groups by two elements

311

Bokowski, J. See A. Altshuler

du Sautoy, M. P. F. Zeta functions of groups and rings: uniformity  1

Fernández-Alcober, G. A. See A. Vera-López

Grunewald, F., J. Mennicke and L. Vaserstein On the groups

SL_{2}(ℤ[x]) and SL_{2}(k[x, y]) 157

Handelman, D. Space-time boundaries for random walks obtained from diffuse measures

107

Jarden, M. and A. Razon Pseudo algebraically closed fields over rings

25

Keller, G. Exponential weak Bernoulli mixing for Collet–Eckmann maps

300

Kwon, K.-H. On the growth of entire functions  409

Marton, K. and P. C. Shields The positive-divergence and blowingup properties

331

Mennicke, J. See F. Grunewald

Mozes, S. On closures of orbits and arithmetic of quaternions  195

Pascuas, D. See C. Berenstein

Razon, A. See M. Jarden

Riley, D. M. and J. F. Semple Completion of restricted Lie algebras 277

Schuchert, P. See A. Altshuler

Semple, J. F. See D. M. Riley

Shelah, S. and L. Soukup On the number of non-isomorphic subgraphs 349 Shields, P. C. See K. Marton

Solomyak, M. Piecewise-polynomial approximation of functions from H^{l}((0, 1)^{d}), 2l = d, and applications to the spectral theory of the Schrödinger operator

253

Soukup, L. See S. Shelah

Vaserstein, L. See F. Grunewald

Vera-López, A. and G. A. Fernández-Alcober The conjugacy vector of a p-group of maximal class 233